Sully Corp. currently has an EPS of $3.90, and the benchmark PE for the company is 38. Earnings are expected to grow at 10 percent per year.

a. What is your estimate of the current stock price? (Round your answer to 2 decimal places. (e.g., 32.16))

Current stock price $

b. What is the target stock price in one year? (Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g., 32.16))

Target stock price $

c. Assuming the company pays no dividends, what is the implied return on the company’s stock over the next year? (Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g., 32.16))

Implied return of stock %
1
Expert's answer
2013-03-13T08:04:08-0400
Sully Corp. currently has an EPS of $3.90, and the benchmark PE for the
company is 38. Earnings are expected to grow at 10 percent per year.
a. What is your estimate of the current stock price? (Round your answer
to 2 decimal places. (e.g., 32.16))
Current stock price = $3.9*38 = $148.2
b. What is the target stock price in one year? (Do not round
intermediate calculations and round your final answer to 2 decimal places.
(e.g., 32.16))
Target stock price = 148.2*1.1 = $163.02
c. Assuming the company pays no dividends, what is the implied return on the
company’s stock over the next year? (Do not round intermediate calculations
and round your final answer to 2 decimal places. (e.g., 32.16))
Implied return of stock = (163.02 - 148.2)/148.2 = 10%